Six copper wires have the lengths and cross-sectional areas listed below. The resistance R of each segment can be calculated using, R = pL/ A

Where p is the resistivity, L is the length of the wire, and A is the wire's cross-sectional area. Rank the resistances of the wires from lowest to highest.



Six copper wires have the lengths and cross-sectio
Answer #1

Given that

        R'=pL/A...... for the first case

        R'=pL/2A=R/2....for the second case

        R'=p(2L)/(A/2) =4pL/A=4R......for the third case

        R'=p(L/2)/(2A) =pL/4A =R/4....for the fourth case

       R'=p(2L)/A =2pL/A =2R....for the first case

From the lowest to highest is

           L/2,2A < L,2A < L,A<2L,A<2L,A/2